OpenStudy (babyslapmafro):

Please help. How do you determine whether the following sequence (click to see) converges or diverges? Thanks.

OpenStudy (babyslapmafro):

\[\left\{( \frac{ n+1 }{ n+3 })^{n} \right\}_{n=1}^{\infty}\]